# and quantum information science

### v2

@@ 9.3 Bounding complexity growth with entanglement
Quantum circuit complexity of a state quantifies the elementary resources necessary to prepare it from a product state. Therefore, complexity also quantifies the resources necessary to transform the state back into a product state. Clearly, therefore, complexity can be viewed as a measure of entanglement, which is ultimately the failure of a state to be of product form. As such, we can compare it to the entanglement entropy, arguably the most prominent measure of entanglement. A high entanglement entropy along a bipartite cut does not imply a high circuit complexity: the maximally entangled state between two systems each consisting of $n$ qubits can be prepared on a quantum computer with a constant depth quantum circuit. Even if we consider the maximum entanglement entropy over all bipartite cuts of the qubits, we find that it saturates trivially at a maximum value of $n$, whereas the circuit complexity can be exponentially large, of order $\Omega(2^{n}/n)$ (compare Section 4.3). On the other hand, we can easily show the existence of quantum circuits with superpolynomial circuit complexity but no entanglement between some parts of the system. For example, simply apply a Haar random unitary to $\log^{1+\varepsilon}(n)$ many qubits and leave all the other qubits untouched. So the circuit complexity and the entanglement entropy are not equivalent; but what can we say about their relationship?
Entanglement growth implies early-time bounds on complexity during chaotic evolution. More precisely, it turns out that entanglement does imply non-trivial circuit lower bounds, but entanglement needs to be present along any bipartite cut [121]. The reason is that the Schmidt rank along any bipartite cut can only be increased by a constant factor. In particular, if a state $|\psi\rangle$ has a Schmidt rank of $\mathrm{SR}_{i}$ for a bipartite cut of all qubits into two sets $A=\{1,\ldots,i\}$ and $B=\{i+1,\ldots,n\}$, then we know that at least $c\log(\mathrm{SR}_{i})$ gates (each acting on a qubit in $A$ and a qubit in $B$) are required to prepare $|\psi\rangle$. Suppose that we only want to lower bound the size of any geometrically local circuit that approximates the state $|\psi\rangle$. A geometrically local circuit can only apply gates to neighbouring qubits $(i,i+1)$. Then, any gate that acts on $A$ and $B$ necessarily acts on $(i,i+1)$. We can thus apply the above argument for all $n$ Schmidt ranks $\mathrm{SR}_{i}$ without overcounting and find:
@@
At best, the left hand side is $\Omega(n^{2})$. This is because each Schmidt rank $\mathrm{SR}_{i}$ is trivially bounded by the state space dimension $2^{n}$. Therefore, it is impossible to prove exponential circuit complexity with this technique. Using similar tools, complexity was lower-bounded in one-dimensional models composed by noninteracting fermions [442]. Reference [443] argued that the area of Hartman-Maldacena surfaces (i.e., codimension-two minimal surfaces homologous to a boundary subregion [245]) lower-bounds the Schmidt rank of the unitary time evolution operator. In turn, the Schmidt rank lower-bounds the depth of any boundary circuit effecting such boundary unitary.
Next, let us mention a relation between circuit complexity and the so-called embezzlement of entanglement. The latter terminology refers to a process in which entanglement is extracted from a resource system by using local unitary operators, with approximately no detectable change to the resource state [444]. The embezzlement error estimates the amount by which this process changed the resource state. References [445, 446] argued that any state of a relativistic QFT could serve as a resource for the process, where the embezzlement error can be made arbitrarily small. Recently, the author of [447] showed that circuit complexity is lower-bounded by a quantity proportional to the inverse of the embezzlement error. In other words, circuit complexity acts as an obstruction to perfect embezzlement.
@@
To introduce the relationship, we first overview relevant background [448, 15]. Consider an $n$-qubit system. The Pauli group consists of $4^{n}$ Pauli strings, each a tensor product of $n$ Pauli operators and/or identity operators. The Clifford group is the Pauli group’s normalizer; each Clifford gate transforms every Pauli string into a Pauli string. The Hadamard, phase, and controlled-NOT gates span the Clifford group. The stabilizer states result from acting on $\lvert 0\rangle^{\otimes n}$ with Clifford gates. According to the Gottesman-Knill theorem, classical computers can efficiently simulate every quantum computation formed only from Clifford gates (such that the quantum computer always remains in the space of stabilizer states) [449]. Consequently, quantum-computational speedups stem from nonstabilizer states. The extra spice added by these states is called magic or nonstabilizerness [450]. This topic has recently advanced rapidly (e.g., [451, 452, 453, 454, 455, 456, 457, 458, 459, 460]). One can quantify magic in various ways, including with the stabilizer Rényi entropy, an entropic measure defined in terms of the Bell basis [453, 460]. The authors of [454] define a magic power that quantifies a unitary’s ability to behave unlike a Clifford gate—to evolve a Pauli string to a nontrivial linear combination of Pauli strings. The magic power lower-bounds the unitary’s circuit complexity [454]. The latter is quantified with Nielsen’s circuit complexity, particularly the cost function (25) evaluated at $p=q_{I}=1$.
@@ 9.4 Bounding complexity via concentration
We briefly sketch a proof of the concentration explained above, where we follow the argument presented in Ref. [461]. Readers that are only interested in how to apply such a statement to circuit complexity can skip this paragraph. $|\psi\rangle$ is the unique ground state of the commuting, gapped and $l$-local Hamiltonian $H=\sum_{i}\mathbbm{1}-U\mathbbm{1}_{i-1}\otimes|0_{i}\rangle\langle 0_{i}|\otimes\mathbbm{1}_{n-i}U^{\dagger}$, where $l\leq 2^{\mathrm{depthof}\,U}$. We can use that Chebyshev polynomials $C_{m}$ (of the first kind) satisfy $C_{m}(x)\leq e^{-m\sqrt{1-h}}$ for $x\in[0,h]$ and $C_{r}(1)=1$. Using that the spectral gap of $H$ is $1$, we find
$\left|\left||\psi\rangle\langle\psi|-C_{m}\left(\mathbbm{1}-\frac{H}{n}\right)\right|\right|_{\infty}\leq e^{-m/\sqrt{n}}.$ (159)
Choosing $m=c^{\prime}n$ for another constant $c^{\prime}>0$ makes this small. However, we notice that the operator $C_{m}\left(\mathbbm{1}-\frac{H}{n}\right)$ is a sum over $lc^{\prime}n$-local terms. In particular, for a string of Hamming weight $<cn$, we have that each bitstring that has overlap with $C_{m}\left(\mathbbm{1}-\frac{H}{n}\right)|x\rangle$ has Hamming weight at most $(c+lc^{\prime})n$. Combining this limitation on the increase in Hamming weight with the approximation (159) yields
@@
This strategy based on concentration in output probability distributions,was used in the proof of Freedman and Hasting’s NLTS (“No Low-Energy Trivial States”) conjecture [462] due to Anshu, Breuckmann and Nirkhe [463]. The NLTS theorem asserts that Hamiltonians exist such that any state with an energy density below some constant threshold is topologically ordered in the sense of Section 9.2. More precisely, there is a family of $k$-local Hamiltonians $H_{n}=\sum_{i=1}^{m}h_{i}$ acting on $n$ qubits with $m=O(n)$ and an $\varepsilon>0$ such that all states $|\psi\rangle$ with $\langle\psi|H|\psi\rangle/n\leq\varepsilon$ are topologically ordered. In particular, this result shows that topological order can persist at constant temperatures.
@@ 10 Epilogue
Sofia: As a matter of fact, Refs. [33, 34, 35, 370] built on [32] to identify a quantitative relationship recently. Section 7.5.1 described this breakthrough: a double-scaled SYK model’s spread complexity is dual to the size of an ERB in two-dimensional theories of gravity. The duality, being exact, is a major achievement.
@@
Sofia: The duality is simplest in two bulk dimensions, and this simplicity facilitated the proof. Yet [35] reinterprets the matching in a language that should enable a generalization to higher dimensions. Additionally, insight into two-dimensional theories has advanced long-standing problems, famously including the black-hole-information paradox [464, 465, 466, 8]. The similar insight here will hopefully enable similar advancements.
@@
Sofia: They do to an extent. Consider an ensemble of theories, such as the SYK model, distinguished by random values of one or more parameters. Variances associated to Krylov and spread complexities diagnose chaos [23, 24, 184]. Furthermore, individual chaotic theories, even if we do not consider an ensemble, have more regular tridiagonal spectra (also called Lanczos spectra) than integrable ones. In fact, there is a precise conjecture for the functional form of the tridiagonal spectral covariances in chaotic theories [24]. Whether and how precisely Krylov complexity distinguishes chaotic from integrable QFTs remains an open question [205, 206], although again the key distinction should lie in the fluctuations of the tridiagonal spectrum of the Hamiltonian. Recent advances appear also in Refs. [207, 208], which leverage von Neumann algebras and algebraic QFTs.
@@
Sofia: Yes, Sec. 5.3 discussed constraints on the complexity metric for finite-dimensional quantum system so that Nielsen complexity shows distinguishing features for chaotic dynamics. Similarly, Sec. 5.4 established physical criteria that lead to long-term Nielsen complexity growth in metrics that penalize complex versus simple generators. Systems that satisfy these criteria, which include some that show spectral features of chaos like level repulsion, show linear growth of Nielsen complexity for exponential time, at least up to possible obstructions arising from the global structure of the unitary group that are hard to study. This long-term complexity growth has its roots in the negative average sectional curvature of the complexity geometry on the unitary group manifold [50, 108, 112, 101]. References [97, 19] discussed various classes of complexity geometries that penalize complex operations as desired. These geometries have a negative average sectional curvature, and are constructed to facilitate analytical computation of geodesics as desired for the study of Nielsen complexity.
@@
Sofia: Certainly—we have known for years that quantum information theory can reveal which part of the bulk encodes which feature of the boundary theory. This research program relates to the claim that slices of AdS geometry naturally arise in tensor networks [239]. The claim inspired notions of complexity discussed in Sec. 8.1. These notions may help us identify holographically dual quantities.
@@
Sofia: Perhaps. Whether de Sitter spacetime has a quantum dual remains an open question. If the answer is yes, then the methods of the AdS/CFT duality may port over to de Sitter spacetime, perhaps with some modifications. If so, we can translate quantum-gravity questions into conceptually simpler questions about non-gravitating quantum systems. For now, people are conjecturing such a translation, as well as conjecturing about duals of complexity [305, 306, 307, 308, 309, 310, 313, 314, 315, 316, 317, 318, 319, 320, 321, 322, 323, 324, 325, 326, 327, 328, 329, 330, 331, 332, 192, 333, 334, 335, 336, 337, 338, 339, 278, 340, 341, 342, 343, 344, 345, 346, 347, 348, 349, 350, 358, 353, 355, 354, 362, 359, 357, 360, 364, 467, 468]. Hopefully, exploring holographic complexity in de Sitter spacetime will indicate which properties the spacetime’s quantum dual should have, if it exists.
Such a development would unlock further opportunities for future work: some researchers expect the quantum dual of de Sitter space to be an open quantum system [192, 193]. If this is the case, we may learn about complexity of dynamics in open quantum systems through de Sitter holography. Several works have already begun unpacking complexity in open quantum systems [469, 470, 471, 429, 472, 473, 474, 475, 476, 477, 478, 479].
Complexio: Also on the topic of holography, according to the review, holographers called for the development of quantum-information-theoretic tools, including a resource theory for quantum complexity (see Sec. 9.1). Quantum information theorists defined the resource theory in [429]. In turn, the resource theory led to new entropic measures of complexity [429, 430]. Can these made-to-order tools now benefit holography?
@@
Sofia: You heard correctly. This skepticism led to the research topic that debuted in [397]: the holographic dictionary’s complexity. According to Sec. 7.7, translating between complexity and volume takes a lot of time unless quantum gravity allows for more computational power than a quantum computer. Computer scientists have proposed that the holographic dictionary might be highly complex. Their argument extends beyond the complexity=volume conjecture to other holographic complexity conjectures. Followup work focused on the duality between area and entanglement entropy [405]. Still, we are not certain for which entries the holographic dictionary is complex. Ramifications will include our ability to study quantum gravity through quantum computing.
@@
Second, we should lower-bound the circuit complexities of states in families that we can describe efficiently. We already have a lower bound $\sim n\,\log(n)$ as discussed in Sec. 9.4. Such an endeavor may encounter obstacles that classical computer scientists face. More precisely, pseudorandom functions could hinder the establishment of such circuit-complexity lower bounds. The reason, roughly, is the following: imagine a quantity that distinguishes low-complexity-functions from high-complexity-functions. This quantity must, itself, be hard to compute. Otherwise, one could easily use the quantity to distinguish pseudorandom functions (low complexity) from random functions (high complexity). The resulting obstacles for lower-bounding classical circuit complexity are called natural proof barriers [70]. We might develop the quantum analog of such natural proof barriers by strengthening a recent construction of pseudorandom unitaries [399].
Third, researchers expect that the low-energy states of some local Hamiltonians have superpolynomial circuit complexities. However, proving an unconditional superpolynomial lower bound would come close to separating the complexity classes $\mathrm{QMA}$ (Quantum Merlin-Arthur) and $\mathrm{QCMA}$ (Quantum-Classical Merlin-Arthur) [480]. QMA contains decision problems that one can verify with a quantum proof; and QCMA, decision problems that a quantum computer can verify with a classical proof. Unconditionally separating QMA and QCMA, one would achieve a monumental breakthrough in theoretical computer science by also separating the classes P and PSPACE (see [481] for details). A more modest goal is to separate QMA from QCMA “relative to an oracle”. An oracle is a black box that solves a powerful problem in one step. Often, unconditional statements about complexity classes can be made if all Turing machines are given access to such an oracle. The problem of separating QMA from QCMA relative to increasingly weaker oracles was the subject of significant efforts in the past years [482, 480, 483, 484, 485, 486, 487, 488].
Let me mention, as an aside, that the mathematics community has studied problems related to circuit optimization. They have focused on the minimum number of steps required to reach all the points on certain group manifolds. We can recognize this number as a type of complexity. Finding gates that minimize this complexity amounts to solving the golden-gates problem [489, 490, 491]. That problem, in turn, relates to the famous mathematical problem of sphere packing. So complexity is really everywhere!

### v3

@@ 9.3 Bounding complexity growth with entanglement
Quantum circuit complexity of a state quantifies the elementary resources necessary to prepare it from a product state. Therefore, complexity also quantifies the resources necessary to transform the state back into a product state. Clearly, therefore, complexity can be viewed as a measure of entanglement, which is ultimately the failure of a state to be of product form. As such, we can compare it to the entanglement entropy, arguably the most prominent measure of entanglement. A high entanglement entropy along a bipartite cut does not imply a high circuit complexity: the maximally entangled state between two systems each consisting of $n$ qubits can be prepared on a quantum computer with a constant depth quantum circuit. Even if we consider the maximum entanglement entropy over all bipartite cuts of the qubits, we find that it saturates trivially at a maximum value of $n$, whereas the circuit complexity can be exponentially large, of order $\Omega(2^{n}/n)$ (compare Section 4.3). On the other hand, we can easily show the existence of quantum circuits with superpolynomial circuit complexity in the system size but no entanglement between some parts of the system. For example, simply apply a Haar random unitary to $\log^{1+\varepsilon}(n)$ many qubits and leave all the other qubits untouched. So the circuit complexity and the entanglement entropy are not equivalent; but what can we say about their relationship?
Entanglement growth implies early-time bounds on complexity during chaotic evolution. More precisely, it turns out that entanglement does imply non-trivial circuit lower bounds, but entanglement needs to be present along any bipartite cut eisert2021entangling. The reason is that the Schmidt rank along any bipartite cut can only be increased by a constant factor. In particular, if a state $|\psi\rangle$ has a Schmidt rank of $\mathrm{SR}_{i}$ for a bipartite cut of all qubits into two sets $A=\{1,\ldots,i\}$ and $B=\{i+1,\ldots,n\}$, then we know that at least $c\log(\mathrm{SR}_{i})$ gates (each acting on a qubit in $A$ and a qubit in $B$) are required to prepare $|\psi\rangle$. Suppose that we only want to lower bound the size of any geometrically local circuit that approximates the state $|\psi\rangle$. A geometrically local circuit can only apply gates to neighbouring qubits $(i,i+1)$. Then, any gate that acts on $A$ and $B$ necessarily acts on $(i,i+1)$. We can thus apply the above argument for all $n$ Schmidt ranks $\mathrm{SR}_{i}$ without overcounting and find:
@@
At best, the left hand side is $\Omega(n^{2})$. This is because each Schmidt rank $\mathrm{SR}_{i}$ is trivially bounded by the state space dimension $2^{n}$. Therefore, it is impossible to prove exponential circuit complexity with this technique. Using similar tools, complexity was lower-bounded in one-dimensional models composed by noninteracting fermions Aravinda:2023pcp. Reference Milekhin:2024mce argued that the area of Hartman-Maldacena surfaces (i.e., codimension-two minimal surfaces homologous to a boundary subregion Hartman:2013aa) lower-bounds the Schmidt rank of the unitary time evolution operator. In turn, the Schmidt rank lower-bounds the depth of any boundary circuit effecting such boundary unitary.
Next, let us mention a relation between circuit complexity and the so-called embezzlement of entanglement. The latter terminology refers to a process in which entanglement is extracted from a resource system by using local unitary operators, with approximately no detectable change to the resource state Hayden_embezz. The embezzlement error estimates the amount by which this process changed the resource state. References vanLuijk:2024fdw, vanLuijk:2024nnx argued that any state of a relativistic QFT could serve as a resource for the process, where the embezzlement error can be made arbitrarily small. Recently, the author of Schwartzman:2024zim showed that circuit complexity is lower-bounded by a quantity proportional to the inverse of the embezzlement error. In other words, circuit complexity acts as an obstruction to perfect embezzlement.
@@
To introduce the relationship, we first overview relevant background 97_Gottesman_PhD, nielsen2010quantum. Consider an $n$-qubit system. The Pauli group consists of $4^{n}$ Pauli strings, each a tensor product of $n$ Pauli operators and/or identity operators. The Clifford group is the Pauli group’s normalizer; each Clifford gate transforms every Pauli string into a Pauli string. The Hadamard, phase, and controlled-NOT gates span the Clifford group. The stabilizer states result from acting on $\lvert 0\rangle^{\otimes n}$ with Clifford gates. According to the Gottesman-Knill theorem, classical computers can efficiently simulate every quantum computation formed only from Clifford gates (such that the quantum computer always remains in the space of stabilizer states) Gottesman:1998hu. Consequently, quantum-computational speedups stem from nonstabilizer states. The extra spice added by these states is called magic or nonstabilizerness Veitch_2014. This topic has recently advanced rapidly (e.g., Liu:2020yso, White:2020zoz, Leone:2021rzd, Bu:2022ozl, Wang:2023uog, Tirrito:2023fnw, Turkeshi:2023lqu, Frau:2024qmf, Cao:2024nrx, Hoshino:2025jko). One can quantify magic in various ways, including with the stabilizer Rényi entropy, an entropic measure defined in terms of the Bell basis Leone:2021rzd, Hoshino:2025jko. The authors of Bu:2022ozl define a magic power that quantifies a unitary’s ability to behave unlike a Clifford gate—to evolve a Pauli string to a nontrivial linear combination of Pauli strings. The magic power lower-bounds the unitary’s circuit complexity Bu:2022ozl. The latter is quantified with Nielsen’s circuit complexity, particularly the cost function (25) evaluated at $p=q_{I}=1$.
@@ 9.4 Bounding complexity via concentration
We briefly sketch a proof of the concentration explained above, where we follow the argument presented in Ref. anshu2023concentration. Readers that are only interested in how to apply such a statement to circuit complexity can skip this paragraph. $|\psi\rangle$ is the unique ground state of the commuting, gapped and $l$-local Hamiltonian $H=\sum_{i}\mathbbm{1}-U\mathbbm{1}_{i-1}\otimes|0_{i}\rangle\langle 0_{i}|\otimes\mathbbm{1}_{n-i}U^{\dagger}$, where $l\leq 2^{\mathrm{depthof}\,U}$. We can use that Chebyshev polynomials $C_{m}$ (of the first kind) satisfy $C_{m}(x)\leq e^{-m\sqrt{1-h}}$ for $x\in[0,h]$ and $C_{r}(1)=1$. Notice that the spectrum of $C_{m}\left(\mathbbm{1}-\frac{H}{n}\right)$ consists of the values $C_{m}(1-\lambda_{i}/n)$, where the $\lambda_{i}$ are the eigenvalues of $H$. Therefore, plugging in that the spectral gap of HH is $1$, we find for the second highest eigenvalue of $C_{m}(1-H/n)$ the following bound:
[…]
Choosing $m=c^{\prime}n$ for another constant $c^{\prime}>0$ makes this small. However, we notice that the operator $C_{m}\left(\mathbbm{1}-\frac{H}{n}\right)$ is a sum over $lc^{\prime}n$-local terms. In particular, for a string of Hamming weight <cn<cn, we have that each bitstring that has overlap with $C_{m}\left(\mathbbm{1}-\frac{H}{n}\right)|x\rangle$ has Hamming weight at most $(c+lc^{\prime})n$. Combining this limitation on the increase in Hamming weight with the approximation (159) yields
@@
This strategy based on concentration in output probability distributions,was used in the proof of Freedman and Hasting’s NLTS (“No Low-Energy Trivial States”) conjecture freedman2013quantum due to Anshu, Breuckmann and Nirkhe anshu2023nlts. The NLTS theorem asserts that Hamiltonians exist such that any state with an energy density below some constant threshold is topologically ordered in the sense of Section 9.2. More precisely, there is a family of $k$-local Hamiltonians $H_{n}=\sum_{i=1}^{m}h_{i}$ acting on $n$ qubits with $m=O(n)$ and an $\varepsilon>0$ such that all states $|\psi\rangle$ with $\langle\psi|H|\psi\rangle/n\leq\varepsilon$ are topologically ordered. In particular, this result shows that topological order can persist at constant temperatures.
@@ 10 Epilogue
Sofia: As a matter of fact, Refs. Lin:2022rbf, Rabinovici:2023yex, Balasubramanian:2024lqk, Heller:2024ldz built on Berkooz:2018qkz to identify a quantitative relationship recently. Section 7.5.1 described this breakthrough: a double-scaled SYK model’s spread complexity is dual to the size of an ERB in two-dimensional theories of gravity. The duality, being exact, is a major achievement.
@@
Sofia: The duality is simplest in two bulk dimensions, and this simplicity facilitated the proof. Yet Balasubramanian:2024lqk reinterprets the matching in a language that should enable a generalization to higher dimensions. Additionally, insight into two-dimensional theories has advanced long-standing problems, famously including the black-hole-information paradox Penington:2019npb, Almheiri:2019psf, Almheiri:2019hni, Almheiri:2020cfm. The similar insight here will hopefully enable similar advancements.
@@
Sofia: They do to an extent. Consider an ensemble of theories, such as the SYK model, distinguished by random values of one or more parameters. Variances associated to Krylov and spread complexities diagnose chaos Balasubramanian:2022dnj, Balasubramanian:2023kwd, Chapman:2024pdw. Furthermore, individual chaotic theories, even if we do not consider an ensemble, have more regular tridiagonal spectra (also called Lanczos spectra) than integrable ones. In fact, there is a precise conjecture for the functional form of the tridiagonal spectral covariances in chaotic theories Balasubramanian:2023kwd. Whether and how precisely Krylov complexity distinguishes chaotic from integrable QFTs remains an open question Avdoshkin:2022xuw, Camargo:2022rnt, although again the key distinction should lie in the fluctuations of the tridiagonal spectrum of the Hamiltonian. Recent advances appear also in Refs. Ouseph:2023juq, Gesteau:2023rrx, which leverage von Neumann algebras and algebraic QFTs.
@@
Sofia: Yes, Sec. 5.3 discussed constraints on the complexity metric for finite-dimensional quantum system so that Nielsen complexity shows distinguishing features for chaotic dynamics. Similarly, Sec. 5.4 established physical criteria that lead to long-term Nielsen complexity growth in metrics that penalize complex versus simple generators. Systems that satisfy these criteria, which include some that show spectral features of chaos like level repulsion, show linear growth of Nielsen complexity for exponential time, at least up to possible obstructions arising from the global structure of the unitary group that are hard to study. This long-term complexity growth has its roots in the negative average sectional curvature of the complexity geometry on the unitary group manifold Brown:2016wib, Brown:2019whu, Balasubramanian:2019wgd, Auzzi:2020idm. References Brown:2021uov, Brown:2022phc discussed various classes of complexity geometries that penalize complex operations as desired. These geometries have a negative average sectional curvature, and are constructed to facilitate analytical computation of geodesics as desired for the study of Nielsen complexity.
@@
Sofia: Certainly—we have known for years that quantum information theory can reveal which part of the bulk encodes which feature of the boundary theory. This research program relates to the claim that slices of AdS geometry naturally arise in tensor networks Swingle:2009bg. The claim inspired notions of complexity discussed in Sec. 8.1. These notions may help us identify holographically dual quantities.
@@
Sofia: Perhaps. Whether de Sitter spacetime has a quantum dual remains an open question. If the answer is yes, then the methods of the AdS/CFT duality may port over to de Sitter spacetime, perhaps with some modifications. If so, we can translate quantum-gravity questions into conceptually simpler questions about non-gravitating quantum systems. For now, people are conjecturing such a translation, as well as conjecturing about duals of complexity Gibbons:1977mu, Bousso:1999dw, Banks:2000fe, Bousso:2000nf, Banks:2001yp, Strominger:2001pn, Banks:2002wr, Leblond:2002ns, Leblond:2002tf, Parikh:2002py, Dyson:2002nt, Dyson:2002pf, Parikh:2004wh, Banks:2005bm, Freivogel:2005qh, Banks:2006rx, Lowe:2010np, Anninos:2011af, Fischetti:2014uxa, Anninos:2017hhn, Anninos:2018svg, Banks:2018ypk, Dong:2018cuv, Lewkowycz:2019xse, Banks:2020zcr, Mirbabayi:2020grb, Anninos:2020cwo, Anninos:2020hfj, Susskind:2021omt, Susskind:2021dfc, Coleman:2021nor, Susskind:2021esx, Shaghoulian:2021cef, Shaghoulian:2022fop, Lin:2022nss, Banihashemi:2022htw, Silverstein:2022dfj, Chandrasekaran:2022cip, Witten:2023xze, Witten:2023qsv, Rahman:2023pgt, Narovlansky:2023lfz, Batra:2024kjl, Verlinde:2024znh, Verlinde:2024zrh, Tietto:2025oxn, Chapman:2021eyy, Jorstad:2022mls, Auzzi:2023qbm, Anegawa:2023wrk, Anegawa:2023dad, Baiguera:2023tpt, Aguilar-Gutierrez:2023zqm, Baiguera:2024xju, Aguilar-Gutierrez:2023pnn, Aguilar-Gutierrez:2024rka, Aguilar-Gutierrez:2024nau. Hopefully, exploring holographic complexity in de Sitter spacetime will indicate which properties the spacetime’s quantum dual should have, if it exists.
Such a development would unlock further opportunities for future work: some researchers expect the quantum dual of de Sitter space to be an open quantum system Anninos:2020cwo, Anninos:2022qgy. If this is the case, we may learn about complexity of dynamics in open quantum systems through de Sitter holography. Several works have already begun unpacking complexity in open quantum systems Bhattacharyya:2021fii, Bhattacharyya:2022rhm, Bhattacharya:2022gbz, YungerHalpern:22Resource, Liu:2022god, Schuster:2022bot, Bhattacharjee:2022lzy, NSSrivatsa:2023pby, Bhattacharyya:2024duw, Ermakov:2024xtn, Schuster:2024jds, Guerra:2025dam.
Complexio: Also on the topic of holography, according to the review, holographers called for the development of quantum-information-theoretic tools, including a resource theory for quantum complexity (see Sec. 9.1). Quantum information theorists defined the resource theory in YungerHalpern:22Resource. In turn, the resource theory led to new entropic measures of complexity YungerHalpern:22Resource, Munson:2024Complexity. Can these made-to-order tools now benefit holography?
@@
Sofia: You heard correctly. This skepticism led to the research topic that debuted in Bouland:2019pvu: the holographic dictionary’s complexity. According to Sec. 7.7, translating between complexity and volume takes a lot of time unless quantum gravity allows for more computational power than a quantum computer. Computer scientists have proposed that the holographic dictionary might be highly complex. Their argument extends beyond the complexity=volume conjecture to other holographic complexity conjectures. Followup work focused on the duality between area and entanglement entropy akers2024holographic. Still, we are not certain for which entries the holographic dictionary is complex. Ramifications will include our ability to study quantum gravity through quantum computing.
@@
Second, we should lower-bound the circuit complexities of states in families that we can describe efficiently. We already have a lower bound $\sim n\,\log(n)$ as discussed in Sec. 9.4. Such an endeavor may encounter obstacles that classical computer scientists face. More precisely, pseudorandom functions could hinder the establishment of such circuit-complexity lower bounds. The reason, roughly, is the following: imagine a quantity that distinguishes low-complexity-functions from high-complexity-functions. This quantity must, itself, be hard to compute. Otherwise, one could easily use the quantity to distinguish pseudorandom functions (low complexity) from random functions (high complexity). The resulting obstacles for lower-bounding classical circuit complexity are called natural proof barriers razborov2004feasible. We might develop the quantum analog of such natural proof barriers by strengthening a recent construction of pseudorandom unitaries ma2024construct.
Third, researchers expect that the low-energy states of some local Hamiltonians have superpolynomial circuit complexities. However, proving an unconditional superpolynomial lower bound would come close to separating the complexity classes $\mathrm{QMA}$ (Quantum Merlin-Arthur) and $\mathrm{QCMA}$ (Quantum-Classical Merlin-Arthur) aaronson2007quantum. QMA contains decision problems that one can verify with a quantum proof; and QCMA, decision problems that a quantum computer can verify with a classical proof. Unconditionally separating QMA and QCMA, one would achieve a monumental breakthrough in theoretical computer science by also separating the classes P and PSPACE (see aaronson2016p for details). A more modest goal is to separate QMA from QCMA “relative to an oracle”. An oracle is a black box that solves a powerful problem in one step. Often, unconditional statements about complexity classes can be made if all Turing machines are given access to such an oracle. The problem of separating QMA from QCMA relative to increasingly weaker oracles was the subject of significant efforts in the past years aharonov2002quantum, aaronson2007quantum, fefferman2015quantum, natarajan2024distribution, li2023classical, liu2023non, liu2024qma, zhandry2024toward.
Let me mention, as an aside, that the mathematics community has studied problems related to circuit optimization. They have focused on the minimum number of steps required to reach all the points on certain group manifolds. We can recognize this number as a type of complexity. Finding gates that minimize this complexity amounts to solving the golden-gates problem PARZANCHEVSKI2018869, sarnak2637letter, evra2022ramanujan. That problem, in turn, relates to the famous mathematical problem of sphere packing. So complexity is really everywhere!